OpenStudy (anonymous):

Find the distance from (-6,3) to the line defined by y=-3x+5

OpenStudy (anonymous):

Are you looking for the shortest difference?

OpenStudy (anonymous):

Umm i think so.. LOL I honestly don't know :/

OpenStudy (anonymous):

It says to express as a radical... if that helps

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Well probably the shortest. Anyway, the shortest distance from (-6, 3) to y = -3x + 5 is a straight 180 degrees line. Meaning, when I connect the point with the y = -3x + 5, the line that connects the two should be straight and not diagonal, that makes sure that this is the shortest distance. So, to find the shortest distance then from (-6, 3) to y = -3x + 5, the two must have the same y value at that point so when I connect the line with (-6, 3), it forms a straight, horizontal line. So lets solve y = -3x + 5 for when y = 3: 3 = -3x + 5 -2 = -3x x = 2/3 --> At (2/3, 3), y = -3x + 5 is the shortest distance away from (-6, 3). To find the length between these 2 points, we just find the difference between the two x-values 2/3 and -6. 2/3 - (-6) = 2/3 + 6 = 20/3 -->Therefore, the shortest distance from (-6, 3) to y = -3x + 5 is 20/3. @ytk080

OpenStudy (anonymous):

Oh okay thanks!